OpenStudy (anonymous):

simplify square root of 75 over 36

OpenStudy (solomonzelman):

is 36 inside the square root?

OpenStudy (anonymous):



OpenStudy (solomonzelman):

\[\sqrt{\frac{75}{36}}=\frac{\sqrt{75}}{\sqrt{36}}=\frac{\sqrt{25 \times 3}}{\sqrt{6 \times 6}}=\frac{5\sqrt{3}}{6}\]